# SuperGaussian: Repurposing Video Models for 3D Super Resolution

### v1

@@ Abstract
We evaluate our proposed SuperGaussian on a variety of 3D inputs, which are diverse both in terms of complexity and representation (e.g., Gaussian Splats or NeRFs), and demonstrate that our simple method significantly improves the fidelity of the final 3D models.
@@
3D super resolution video upsampling category-agnostic 3D generation gaussian splatting1 Introduction
@@
We address this challenge with a simple, modular, and generic approach that can be integrated into existing workflows. We proceed as follows: First, we render a video of the scene from the coarse 3D input given sampled view trajectories; second, we upsample the rendered video using a pretrained video-based upsampler that is optionally fine-tuned to handle artifacts of the input modality. For 3D consolidation, we adopt Gaussian Splatting [27] as our output representation. Being an object-centric representation, Gaussian splats are ideally suited for encoding individual objects and are capable of capturing local details. Gaussian splats also strike a good balance between simplicity, fidelity of encoded models, and efficiency of rendering. Fig. 1 shows that our proposed video-based 3D super-resolution framework can be successfully adapted to upscale diverse sets of coarse 3D modalities.
@@ 2 Related Work
Super-Resolution for Video. Following the success of learning-based single-image super-resolution methods, there have been various efforts to extend this success to the video domain. Several strategies have been explored to make this extension. While some work [33] introduces temporal aggregation modules to aggregate information across time, recurrent neural networks [60, 8] with uni- and bi-directional information sharing have also been a choice of architecture. Deformable convolutions [52, 48] have been utilized to align per-frame features both for video upsampling and restoration tasks. More recently, diffusion-based video upsamplers have been proposed by injecting temporal layers into base diffusion model [69]. In our work, we leverage VideoGigaGAN [57], a state-of-the-art generative video upsampler. However, our method is agnostic to the video model we use and can further be improved in the future with advances in video upsampling techniques, e.g., diffusion-based video upsampler.
Image-Based Super-Resolution for 3D. Several approaches have recently aimed at increasing the resolution of implicit radiance fields, in particular Neural Radiance Fields (NeRFs). This is an important and practical application for both capture and synthesis scenarios where higher detail and quality are desirable. Starting with a low-res radiance field, at each iteration, this thread of work [19, 23, 49] renders images from the current representation. Next, a pre-trained image up-sampler is used to upsample the renderings. Then the gradient flow coming from the difference between the current renderings and upsampled images, is used to optimize the 3D representation. Similarly, CROP [61] also leverages a single image upsampler but introduces an uncertainty volume to improve the multi-view consistency. Unlike these methods, however, we adopt a video upsampler as opposed to a single-image upsampling network. Thanks to better handling of temporal consistency, we show that the use of the video upsampler improves the visual quality of the final results. Moreover, unlike previous work which mainly focuses on upsampling NeRF representations, SuperGaussian is a generic approach, which we demonstrate on various 3D input types.
@@
Feed-forward 3D Super Resolution. Several 3D generation and reconstruction methods have proposed direct 3D superresolution modules. Smith et al. [47] represent a 3D shape as multi-view orthographic depth maps which are then upsampled. The recent diffusion-based 3D generation method LASDiffusion [68] presents a method to predict signed distance fields from an occupancy map to improve the resolution of the 3D shapes. Both methods, however, focus only on geometry refinement while our work jointly performs appearance and geometry upsampling. Rodin [51], a triplane-based diffusion method, presents a triplane upsampling strategy. This approach is applicable only to the specific triplane features learned. Finally, our approach does not need large-scale 3D datasets and instead leverages existing video models trained on video data.
3D Implicit Representations. Recently, neural implicit representations such as NeRF [38], VolSDF [59], and NeuS [50] have been popularized as an efficient and compact representation for 3D shapes. Efficiently capturing high-resolution details, however, still remains a challenge for such representations due to the low-dimensional volumetric structure and the computationally intensive volumetric rendering process. Recently, Gaussian Splatting [27] has been proposed as an alternative that provides an object-centric representation with fast rendering capabilities. The scene is represented by a large number of 3D Gaussian blobs whose properties are optimized through differentiable rendering. In our work, we adapt Gaussian splats as the final 3D representation output due to the efficient optimization and rendering capabilities they offer.
@@ 4.1 Setup
| image upsampler | 1.14 | |||
| video upsampler | 0.1511 | 6.04 | 14.61 | 1.19 |
@@ 4.2 Comparison Studies
| Instruct-G2G [20] | 1.06 | |||
| Super-NeRF [19] | 1.03 | |||
| Pre-hoc image | 0.99 | |||
| Ours | 1.08 |
@@
| FastSR-NeRF [32] | |||
| NeRF-SR [49] | |||
| Ours |

### v4

@@ Abstract
We evaluate our proposed SuperGaussian on a variety of 3D inputs, which are diverse both in terms of complexity and representation (e.g., Gaussian Splats or NeRFs), and demonstrate that our simple method significantly improves the fidelity of current generative 3D models.
@@
3D super resolution video upsampling category-agnostic 3D generation Gaussian splatting 3D-consistent1 Introduction
@@
We address this challenge with a simple, modular, and generic approach that can be integrated into existing workflows. We proceed as follows: First, we render a video of the scene from the coarse 3D input given sampled view trajectories; second, we upsample the rendered video using a pretrained video-based upsampler that is optionally fine-tuned to handle domain-specific artifacts of the input modality. For 3D consolidation, we adopt Gaussian Splatting [27] as our output representation. Being an object-centric representation, Gaussian splats are ideally suited for encoding individual objects and are capable of capturing local details. Gaussian splats also strike a good balance between simplicity, fidelity of encoded models, and efficiency of rendering. Fig. 1 shows that our proposed video-based 3D super-resolution framework can be successfully adapted to upscale diverse sets of coarse 3D modalities.
@@ 2 Related Work
Video Super-Resolution. Following the success of learning-based single-image super-resolution methods, there have been various efforts to extend this success to the video domain. Several strategies have been explored to make this extension. While some work [33] introduces temporal aggregation modules to aggregate information across time, recurrent neural networks [60, 8] with uni- and bi-directional information sharing have also been a choice of architecture. Deformable convolutions [52, 48] have been utilized to align per-frame features both for video upsampling and restoration tasks. More recently, diffusion-based video upsamplers have been proposed by injecting temporal layers into base diffusion model [69]. In our work, we leverage VideoGigaGAN [57], a state-of-the-art generative video upsampler. However, our method is agnostic to the video model we use and can further be improved in the future with advances in video upsampling techniques, e.g., diffusion-based video upsampler.
Image-Based 3D Super-Resolution. Several approaches have recently aimed at increasing the resolution of implicit radiance fields, in particular Neural Radiance Fields (NeRFs). This is an important and practical application for both capture and synthesis scenarios where higher detail and quality are desirable. Starting with a low-res radiance field, at each iteration, this thread of work [19, 23, 49] renders images from the current representation. Next, a pre-trained image up-sampler is used to upsample the renderings. Then the gradient flow coming from the difference between the current renderings and upsampled images, is used to optimize the 3D representation. Similarly, CROP [61] also leverages a single image upsampler but introduces an uncertainty volume to improve the multi-view consistency. Unlike these methods, however, we adopt a video upsampler instead of a single-image upsampling network. Thanks to better handling of temporal consistency, we show that using the video upsampler improves the visual quality of the final results. Moreover, unlike previous work mainly focusing on upsampling NeRF representations, SuperGaussian is a generic approach, which we demonstrate on various 3D input types.
@@
Feed-forward 3D Super-Resolution. Several 3D generation and reconstruction methods have proposed direct 3D superresolution modules. Smith et al. [47] represent a 3D shape as multi-view orthographic depth maps which are then upsampled. The recent diffusion-based 3D generation method LASDiffusion [68] presents a method to predict signed distance fields from an occupancy map to improve the resolution of the 3D shapes. Both methods, however, focus only on geometry refinement while our work jointly performs appearance and geometry upsampling. Rodin [51], a triplane-based diffusion method, presents a triplane upsampling strategy. This approach is applicable only to the specific triplane features learned. Finally, our approach does not need large-scale 3D datasets and instead leverages existing video models trained on video data.
Deep Implicit Representations. Recently, neural implicit representations such as NeRF [38], VolSDF [59], and NeuS [50] have been popularized as an efficient and compact representation for 3D shapes. Efficiently capturing high-resolution details, however, still remains a challenge for such representations due to the low-dimensional volumetric structure and the computationally intensive volumetric rendering process. Recently, Gaussian Splatting [27] has been proposed as an alternative that provides an object-centric representation with fast rendering capabilities. The scene is represented by a large number of 3D Gaussian blobs whose properties are optimized through differentiable rendering. In our work, we adapt Gaussian splats as the final 3D representation output due to the efficient optimization and rendering capabilities they offer.
@@ 4.1 Setup
| image upsampler | 0.1533 | 7.06 | 14.67 | 11.67 1.14 |
| video upsampler | 0.1511 | 6.04 | 14.61 | 12.45 1.19 |
@@ 4.2 Comparison Studies
| Instruct-G2G [20] | 0.1867 | 8.33 | 32.56 | 10.52 1.06 |
| Super-NeRF [19] | 0.2204 | 8.84 | 37.54 | 10.40 1.03 |
| Pre-hoc image | 0.1524 | 7.65 | 27.04 | 11.27 0.99 |
| Ours | 0.1290 | 6.80 | 24.32 | 11.69 1.08 |
@@
| FastSR-NeRF [32] | 0.075 | 30.47 | 0.944 |
| NeRF-SR [49] | 0.076 | 28.46 | 0.921 |
| Ours | 0.067 | 28.44 | 0.923 |
@@ 5 Conclusion
6 Acknowledgement
We would like to thank Yiran Xu, Difan Liu, and Taesung Park for their discussions related to VideoGigaGAN [57]. Their support has been essential to the development and success of SuperGaussian. We also thank Nathan Carr for providing Gaussian Splat examples.